OpenStudy (anonymous):

2/5(x+1)=g. solve for x.

OpenStudy (anonymous):

You would first multiply both sides by 5. So, 2(x+1)=5g

hero (hero):

Hint: \(\large\frac{2(x+1)}{5} = g\)

OpenStudy (anonymous):

she cant multiply by 5 since its a fraction. she has to mult the parenthesis by 2/5 instead

OpenStudy (anonymous):

Okay.. I got \[\frac{ 5g-2 }{ 2}\] Is that wrong?

hero (hero):

Yup, that's right. And @magepker728, \(\frac{2}{5}(x+1) = \large\frac{2(x+1)}{5}\) . Do you agree?

hero (hero):

It's easier to set up the right side then multiply both sides by 5 and continue solving from there. Some students do not feel comfortable multiplying by fractional inverses.

hero (hero):

@Victorialee seems to agree with me.

OpenStudy (anonymous):

I do! That is much easier that way.

OpenStudy (phi):

\[ \frac{5g-2}{2} \] is correct. But if it is multiple choice, be sure to know that you could also see it as \[ \frac{5g}{2} -1 \] or \[ \frac{5}{2}g-1 \]

OpenStudy (anonymous):

a thumbs up for hero, as for me fraction is the way to go :D

hero (hero):

@magepker728, I take it that you're a student who is not a big fan of alternative methods. You tend to solve math problems exactly the way you were shown in school, correct?

OpenStudy (anonymous):

well I'm majored in computer science first year, and i tend to challenge myself specially with anything with fractions since not to many people love fraction. sorry if there is any error English isn't my first language.